The base of the triangle exceeds the height by 8 meters. If the area is 42 meters, find the length of the base and the height of the triangle.
Found 2 solutions by TimothyLamb, richwmiller:
Answer by TimothyLamb(4379)

b = h + 8
---
a = (1/2)bh
---
a = (1/2)bh
20 = (1/2)(h + 8)h
20 = (1/2)(hh + 8h)
(1/2)hh + 4h - 20 = 0
(1/2)h^2 + 4h - 20 = 0
---
the above quadratic equation is in standard form, with a=1/2, b=4, and c=-20
---
to solve the quadratic equation, plug this:
0.5 4 -20
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
h = 3.48331477
h = -11.4833148
---
the negative root doesn't make sense for a length, so use the positive root:
---
Answer:
b = 11.483 m
h = 3.483 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by richwmiller(17219)

1/2 h*h=a
1/2b*h=42
b*h=84
b=h+8
h=38
and b =46